Staty Geom has a total investment at $24000 at two interest rates: 7% and 9%. If she invested thrice as much amount at 9% than 7%. How much interest will the investment gain in a year?
Answer by richwmiller(17219)

We know that the investment for the two accounts is $24000
We know that the account at 9% has 3 times the amount at 7%
x+y=24000
x=3y
3y+y=24000
4y=24000
y=$6000 at 7%
x=$18000 at 9%
Your total for x is $18000.00
Your total for y is $ 6000.00
Total invested $18000.00 + $ 6000.00=$24000.00
We check
0.09*18000.00+0.07* 6000.00=24000
1620.00+ 420.00=24000
2040=24000
Since this statement is TRUE and neither x nor y is negative then all is well.
codeintri